I have two questions that I need help answering?
1.When simplifying the rational expression, (x+8)/(x +2) explain why it is improper to cancel out the x’s. State a general rule for canceling factors in a rational expression and give an example of how this rule would be used.
2.How is a rational expression different from a rational equation? Explain, in your own words, the process for solving a rational equation. Why is it important to find the LCD when solving a rational equation? Is it possible to solve a rational equation without finding the LCD? Give an example.

Answer by stanbon(75887)

1.When simplifying the rational expression, (x+8)/(x +2) explain why is it improper to cancel out the x’s.
The x's are not factors of the numerator or of the denominator; they are terms.
------
State a general rule for canceling factors in a rational expression and give an example of how this rule would be used.
---
Only factors that are factors of the whole numerator and the whole denominator
may be canceled.
-------
[x(x+2)]/[2x(x-3)] = (x+2)/[2(x-3)]
========================================
2.How is a rational expression different from a rational equation? Explain, in your own words, the process for solving a rational equation.
The equation contains a rational expression AND an equal sign.
---------------------------
Why is it important to find the LCD when solving a rational equation?
---
Only fractions with the same denominator can be added or subtracted.
---------------------------
Is it possible to solve a rational equation without finding the LCD? Give an example.
Yes:
Ex: (x+2)/(x-3) = 6
Cross-multiply to get:
x+2 = 6x-18
5x = 20
x = 4
-----
Cheers,
Stan H.
=============